OpenStudy (anonymous):

Determine if the lines are parallel, perpendicular , same line, or none of the above y = -3/5x+2 and y= 5/3x+2

OpenStudy (anonymous):

parallel line have the same slope. perpendicular lines have slopes m1 and m2, where m1m2=-1. Here the slopes are (-3/5)(5/3)=-1 so the lines are perpendicular

OpenStudy (anonymous):

thanks!!! Will you help me with one more? :)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

okay so its the same deal determining weather lines are parallel perpendicular or the same y = -3/5x+2 and 3x+5y = 20

OpenStudy (anonymous):

re-write the second line as y=(-3/5)+4 compare with first line y=(-3/5)x+2

OpenStudy (anonymous):

are the slope the same or different?

OpenStudy (anonymous):

same?

OpenStudy (anonymous):

yes. the slopes are both -3/5

OpenStudy (anonymous):

now...are they the same line? another way to say this would be: are the two equations identical?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the two equations are the same? are you sure?

OpenStudy (anonymous):

no they are not the same they have the same slope

OpenStudy (anonymous):

(-3/5)x+4 (-3/5)x+2

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so they are different lines.

OpenStudy (anonymous):

but they have the same slope so, this means...?

OpenStudy (anonymous):

they are parallel?

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

Thank you!!!!!!

OpenStudy (anonymous):

welcome